OpenStudy (anonymous):

Need a bit of guidance, the problem is "find the bsolute extreme of f(x)=3x^(2/3)-2x on [-1,2]. Can anyone help?

OpenStudy (anonymous):

f'(x) = (2/3)*3(x^(-1/3)) - 2 = 0 (x^(1/3)) = 1 x = 1 look the derivative cant have x=0 so we also have to check this point i dont know if its minimum or maximum but since we need the absolute: f(0) = 0 f(-1) = 5 f(2) = 0.7622 f(1) = 1 so (-1,5) (0,0)

OpenStudy (anonymous):

Thank you for the help, but I don't quite follow how you got your values for f(x), I got some different numbers, could you show me how you got yours? Thank you again for the help

OpenStudy (anonymous):

x=1 from differentiation.. f'(x) = 0 x=0 also from differentiation where the derivative there is discontinuous.. so i check what is going on with 0. x=-1 x=2 bounds .. function gets minimum or maximum on its bounds..

OpenStudy (anonymous):

Okay, I understand this, but I don't follow how you got f(-1) = 5, I got -1, so my bounds would be (if I am thinking correctly) f(1)=1 and f(-1)=-1

OpenStudy (anonymous):

(-1)^(2/3) = ((-1)^2)^(1/3) = 1 3*1 - (2)*(-1) = 3 + 2 = 5

OpenStudy (anonymous):

3*1 - (2)*(-1) = 3 + 2 = 5, I keep getting 3(-1^2/3) is equal to negative 3, isn't -1^2/3 just equal to -1? thank you again for your help

OpenStudy (anonymous):

you have to do (-1)^(2/3)

OpenStudy (anonymous):

you can either : ((-1)^2)^(1/3) = (1)^(1/3) = 1 or : ((-1)^(1/3))^2 = (-1)^2 = 1

OpenStudy (anonymous):

derivative of \[3 x^{2/3}-2 x \]is\[\frac{2}{\sqrt[3]{x}}-2 \]Combine the terms,set the result to zero and solve for x\[-\frac{2 \left(\sqrt[3]{x}-1\right)}{\sqrt[3]{x}}=0\]\[\sqrt[3]{x}-1=0 \]\[\left(\sqrt[3]{x}\right)^3=1^3 \]\[x=1 \]Plot attached.

OpenStudy (anonymous):

robtobey - you suggest that x^(2/3) is not defined for x<0 ?

OpenStudy (anonymous):

Made a mistake. 3x^(2/3) - 2x on the plot is defined at x = 0.